OpenStudy (anonymous):

I get how to find the terms with the equation but not how to find equation with the terms PLease help :) Find an equation for the nth term of the arithmetic sequence. -20, -16, -12, -8, ...

OpenStudy (***[isuru]***):

can ya be more specific?? Do ya want to build /prove the formula \[T _{n} = a + (n-1)d\] ??

OpenStudy (anonymous):

I think so bc these are my choices and I know its not a an = -20 x 4(n - 1) an = -20 + 4(n - 1) an = -20 + 4(n + 2) an = -20 + 4(n + 1)

OpenStudy (***[isuru]***):

so... this is a multiple choice problem... which means u don;t have to prove that formula... But do u know about that expression ; i mean the equation to find the nth term ??

OpenStudy (anonymous):

all I know is that the first letter is going to be -20 bc that is a1 and the second is adding 4 so the equation will bc an=-20+4 but I don't know what the (n-1) stuff means or to figure out which one I need

OpenStudy (***[isuru]***):

Okieee.... this is an arithmetic progression ... wt is the special feature of an arithmetic progression.... i mean how do u find which one is arithmetic and which one is not if u have 2 progressions ?

OpenStudy (anonymous):

im just now starting to teach myself I know its arethmitic when there is a pattern like when youre multiplying the number by 2 to get the next answer I haven't done geometric yet Im just trying to figure out how to find the last bit of that equation and ill most likely learn the other stuff in my next lesson

OpenStudy (***[isuru]***):

in an arithmetic progression ya keep adding a constant value to the previous term to find the next term... if the first term of an arithmetic progression is "x" and the constant value or the difference between 2 successive terms is "y" 1 st term x + (0xy) = x 2nd term x + (1 x y) = x + y 3rd term x + ( 2 x Y ) = x + 2y 4th term x + (3 x y) = x + 3y did u get it so far ???

OpenStudy (anonymous):

uhm is this how you get this part (n - 1) (n + 2) (n + 1)

OpenStudy (***[isuru]***):

yep!!! did u get it??

OpenStudy (anonymous):

I thought the constant value was an = -20 + 4(n + 2) the +4 because that is what we are adding to a1 to get the next term?

OpenStudy (***[isuru]***):

yeah!! .... we add that value to get the next value...

OpenStudy (***[isuru]***):

but...

OpenStudy (***[isuru]***):

what should we do if we have to find 3rd term from the 1st term.. i.e if the 1st term is 4 and the difference is 2... we will add 2 to 4 in order to find the 2nd term.. which is 4 + 2 = 6 but to find third term... we will have to add that value again to the 2nd term which is 6 + 2 = 8 but we can write this as... 2nd term + 2 =8 ( 4 + 2) + 2 = 8 so what we have done is 4 + (2x2) =8 adding the multiplied difference to the first term... did u get it ?

OpenStudy (anonymous):

so is it this one then ?an = -20 + 4(n + 2)

OpenStudy (***[isuru]***):

no.....

OpenStudy (anonymous):

dang this last part of the equation is dumb let me try reading over your stuff again

OpenStudy (***[isuru]***):

so far... the example we write can be summarize as follow nth term = firs term +[ difference x something ] I think wt troubles u is this "something" part am I right ?

OpenStudy (anonymous):

yea I just don't get how to get the (n-whatever number goes here part)

OpenStudy (anonymous):

but I get how to get the rest of the equation up to the part in the parenthesis

OpenStudy (***[isuru]***):

so.... lets take the example we right.... if the first term of an arithmetic progression is "x" and the constant value or the difference between 2 successive terms is "y" 1 st term x + (0xy) = x 2nd term x + (1 x y) = x + y 3rd term x + ( 2 x Y ) = x + 2y 4th term x + (3 x y) = x + 3y Now observe carefully how we got the 4th term... remember it should be in the form of ... nth term = firs term +[ difference x something ] 4th term x + (3 x y) = x + 3y x + 3y = x + (3 x y) here... nth term is 4 which is x + 3y so we know what is represent by nth term here and the first term is x.... and here comes the part u have problem with... [ difference x something ] = 3y we know the difference is "Y" so... the part "something" is equal to = 3 the term (n-1) is the key to find this... so whatever that part is we should get the answer 3 in this instance... this term we consider here is 4th term and the value of that "something" is 3 so.. we can write "something" = 4th term -1 do u agree with me ?

OpenStudy (***[isuru]***):

in the same way if we move to the 3rd term... x + ( 2 x Y ) = x + 2y here the [difference x "something"] part is (2 x y) so.. the something is 2 which means it can be written as ( 3 - 1 )

OpenStudy (***[isuru]***):

so ... if u want to find the n th term the "something" part for n th term can be written as (n - 1) so the form can be rearrange as.. nth term = firs term +[ difference x something ] = first term + [difference x (n-1) ] did u get it now ???

OpenStudy (anonymous):

so its always going to be written like that ? (n-1)?

OpenStudy (***[isuru]***):

yeahhhh!!!!

OpenStudy (anonymous):

wow awesome thank you ok I have another one same subject different question. if you cant help Its ok jw if you can :)?

OpenStudy (***[isuru]***):

u r welcome!!

OpenStudy (anonymous):

Find an equation for the nth term of the arithmetic sequence. a19 = -58, a21 = -164

OpenStudy (anonymous):

an = 896 - 53(n - 2) an = 896 - 53(n - 1) an = 896 + 53(n + 1) an = 896 - 53(n + 1)

OpenStudy (anonymous):

wait is it b since it has (n-1)?

OpenStudy (***[isuru]***):

yep!!!

OpenStudy (anonymous):

awesome thank you sir!

OpenStudy (***[isuru]***):

no .... wait!! gimme a minute... It could be difference this time

OpenStudy (anonymous):

okie

OpenStudy (***[isuru]***):

first what is the first term of this progression ...?

OpenStudy (***[isuru]***):

ya don't have to calculate it...'cause every answer has the same first term 896

OpenStudy (anonymous):

yea I was just gunna say that lol

OpenStudy (anonymous):

and obvi youre subtracting 53

OpenStudy (anonymous):

so its for sure not c

OpenStudy (***[isuru]***):

yeah!! the difference is -53

OpenStudy (anonymous):

then we get to the dumb part -_-

OpenStudy (***[isuru]***):

and u r correct it is indeed B

OpenStudy (anonymous):

ok and for this last problem I have, a word problem is it the same thing just making the equation? A certain species of tree grows an average of 0.5 cm per week. Write an equation for the sequence that represents the weekly height of this tree in centimeters if the measurements begin when the tree is 800 centimeters tall.

OpenStudy (***[isuru]***):

nth term = first term + [ differance x (n-1)] = 896 -56(n-1)

OpenStudy (anonymous):

so its an=800+.05(n-1)?

OpenStudy (***[isuru]***):

u got it bro...!!!

OpenStudy (anonymous):

sweet thank you so much :)

OpenStudy (***[isuru]***):

u r welcome!!